OpenStudy (anonymous):

Write function in vertex form Y=X^2+6x+2

OpenStudy (anonymous):

What are your choices?

OpenStudy (anonymous):

start by completing the square

OpenStudy (anonymous):

y-2 = x^2 + 6x

OpenStudy (anonymous):

y-2+9 = x^2 +6x + 9

OpenStudy (anonymous):

y+7 = (x+3)^2 y = (x+3)^2 - 7